Let a,b,x be three real numbers with a>b and x>0. Which of the following statements is correct?

A. Xa>Xb if a,b>1 and for every x>0

B. Xa<Xb if X is an element of (0,1)

C.Xa<Xb if a, b>0 and for every x>1

D.Xa>Xb if a,b x>0


1
Expert's answer
2021-07-19T05:53:12-0400

A.

"x^a>x^b, a>b>1,x>0"

 False.

Counterexample


"x=\dfrac{1}{2}>0, a=3>1, b=2>1, 3>2"

"x^a=(\dfrac{1}{2})^3=\dfrac{1}{8}, x^b=(\dfrac{1}{2})^2=\dfrac{1}{4}"


"\dfrac{1}{8}<\dfrac{1}{4}=>x^a<x^b"



B.

"x^a<x^b, a>b,0<x<1"

 True.

The function "f(t)=x^t" is decreasing, if "0<x<1." Then


"x^a<x^b, \text{if }a>b"

C.

"x^a<x^b, a>b>0,x>1"

 False.

Counterexample


"x=2>1, a=3>0, b=2>0, 2>1"

"x^a=(2)^3=8, x^b=(2)^2=4""8>4=>x^a>x^b"


D.

"x^a>x^b, a>b>0,x>0"

 False.

Counterexample


"x=\dfrac{1}{2}>0, a=3>0, b=2>0, 3>2"

"x^a=(\dfrac{1}{2})^3=\dfrac{1}{8}, x^b=(\dfrac{1}{2})^2=\dfrac{1}{4}"


"\dfrac{1}{8}<\dfrac{1}{4}=>x^a<x^b"